OpenStudy (anonymous):

What are the next 3 terms of the following sequence? -3, -8, -13

OpenStudy (anonymous):

a=-3 ,d=-5 tn=a+(n-1)d t4=-3+3*-5=-18 t5=-3+4*-5=-23 t6=-3+5*-5=-28

OpenStudy (anonymous):

:) thank you!